How many gallons of a 50% antifreeze solution must be mixed with 70 gallons of 10% antifreeze to get a mixture that is 40% antifreeze?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

How many gallons of a 50% antifreeze solution must be mixed with 70 gallons of 10% antifreeze to get a mixture that is 40% antifreeze?
----------------
active + active = active
0.5x + 0.10*70 = 0.4(x+70)
---------------------------
50x + 10*70 = 40x + 40*70
10x = 30*70
----
x = 210 gallons (amt of 50% antifreeze needed)
-------------
Cheers,
Stan H.
----------------

Answer by Alan3354(69443)

How many gallons of a 50% antifreeze solution must be mixed with 70 gallons of 10% antifreeze to get a mixture that is 40% antifreeze?
--------
Add G gallons.
70*10 + G*50 = (70+G)*40